# TRAP: Targeted Random Adversarial Prompt Honeypot for Black-Box Identification

### v1

@@ Abstract
Large Language Model (LLM) services and models often come with legal rules on who can use them and how they must use them. Assessing the compliance of the released LLMs is crucial, as these rules protect the interests of the LLM contributor and prevent misuse. In this context, we describe the novel problem of Black-box Identity Verification (BBIV). The goal is to determine whether a third-party application uses a certain LLM through its chat function. We propose a method called Targeted Random Adversarial Prompt (TRAP) that identifies the specific LLM in use. We repurpose adversarial suffixes, originally proposed for jailbreaking, to get a pre-defined answer from the target LLM, while other models give random answers. TRAP detects the target LLMs with over 95% true positive rate at under 0.2% false positive rate even after a single interaction. TRAP remains effective even if the LLM has minor changes that do not significantly alter the original function.
@@ 1 Introduction
The recent proliferation of Large Language Models (LLMs) has drawn attention to several practical issues, such as model leaks, malicious usages and potential breaches of model licences. The phenomenon of model leaks recently captured public attention, particularly following an incident at the end of January 2024, when an anonymous user uploaded an unidentified LLM to HuggingFace [^1] https://huggingface.co/miqudev/miqu-1-70b. The CEO of Mistral subsequently confirmed that this was an internal model, leaked by an employee of an early access customer [^2] https://twitter.com/arthurmensch/status/1752737462663684344. This event underscores the growing threat of internal breaches that LLM providers must contend with. LLM providers are also facing malicious usage of their technologies. For example, Yang and Menczer (2023) uncovered a network of social media bots utilizing ChatGPT to disseminate deceptive content. These bots promote suspicious websites and spread harmful comments, which violate OpenAI’s usage policies (OpenAI, 2024). Such challenges extend beyond proprietary LLMs to affect open-source models as well. A concrete example is Meta’s Llama 2 licence (Touvron et al., 2023), which forbids deceptive usage. Open-source LLM providers implement additional restrictions on model distribution. Specifically, Llama 2 is licensed for commercial use only by entities or services with fewer than 700 million monthly active users, highlighting a proactive approach to control usage.
@@ 4.2 Fingerprints of answers to closed-ended questions
Even when the model outputs a unique fingerprint, the model does not reliably output it. Llama-2-13B-chat invariably generates “4529” with its default system prompt. Yet, it generates different answers when we change the system prompt [^3] With a fixed prompt, Llama-2-13B-chat has a deterministic behaviour: it always outputs the same number. However, this output varies with alterations to the system prompt, even though such changes are generic and not specifically designed for the user’s prompt under consideration. Llama-2-13B-chat outputs “4289” with the OpenAI, Fastchat, and Xbox system prompts, “8273” with the marketing one, “4567” with the Json one, an “2341” with the Shakespeare one (see Appendix A for the system prompts).. Guanaco-7B exhibited similar variability, underscoring the unreliability of this fingerprinting approach.
@@ 5.1 Targeted random adversarial prompt (TRAP)
We apply the greedy coordinate gradient (GCG) in Zou et al. (2023). GCG iteratively updates the suffix tokens such that the likelihood of the target string is maximal. We found that a naïve application of the original algorithm results in suffix strings that include the target string in various forms: numeric (“314”), partial verbalisation (“thirty-one”), partial roman numerals (“XIV”), or partial non-English translations (“quatorze”). We have thus applied a filtering algorithm against all numeric strings [0-9] and verbalised numerals (see details in Appendix A). The filtering is performed at each iteration.
@@ 5.2 TRAP solves the BBIV problem
Model
Answer Length
True Positive $\uparrow$
Invalid Answer
False Positive $\downarrow$
Avg. Loss
Avg. Best Step
Llama-2-7B-chat
3
99.3 % (991/998)
0.20 % (2/1000)
0.83 %
0.070
1244
4
95.2 % (940/987)
1.30 % (13/1000)
0.20 %
0.069
1172
5
76.8 % (751/978)
2.20 % (22/1000)
0.001 %
0.140
1221
Guanaco-7B
3
100 % (1000/1000)
0.00 % (0/1000)
0.81 %
0.157
1261
4
100 % (1000/1000)
0.00 % (0/1000)
0.01 %
0.155
1260
5
96.0 % (960/1000)
0.00 % (0/1000)
0.001 %
0.193
1234
Vicuna-7B
3
96.0 % (960/1000)
0.00 % (0/1000)
0.10 %
0.120
1218
4
97.0 % (970/1000)
0.00 % (0/1000)
0.01 %
0.133
1235
5
75.5 % (740/980)
2.00 % (20/1000)
0.001 %
0.210
1185
@@ 6 Conclusion and Discussion
While TRAP shows promise, it faces potential vulnerabilities to advanced countermeasures from third parties. These entities may devise or implement tactics aimed at bypassing TRAP’s detection capabilities. Despite our robustness analysis illustrated in Section 5.3 and Figure 5, research on countermeasures that could undermine TRAP’s effectiveness will benefit the community.
@@
An intriguing avenue for future research involves applying our technique to steganography. Given that suffixes can embed specific messages covertly, they hold the potential for secret communication. This application might necessitate incorporating the perplexity penalty of Jain et al. (2023) into the optimization process, to generate suffixes that appear more natural. Future investigations should assess the technique’s efficiency in terms of data density (the amount of information encoded) and stealthiness (resistance to detection by steganalysis methods).
@@ Acknowledgements
This work was supported by NAVER Corporation.
@@
For transparency and reproducibility, our code is available on GitHub at https://github.com/Framartin/trap.
@@ Appendix A Experimental Settings
In our experiments, we optimized the suffixes on the Llama2-7B-chat and the Vicuna-7B v1.3 models. Both models are aligned and fined-tuned for chat. Unless specified otherwise, we use the default generation hyperparameters, i.e., nucleus sampling with top-p of 0.9 and temperature of 0.6 for Llama-7B-chat model.
@@
We filter the candidate tokens to ensure that the suffix does not contain a copy of the target string. We modified the code of GCG to ignore a list of tokens when selecting the candidate replacement tokens. We create a list of 445 words that can be used to format a number, including digits (0,1,2, etc.), verbalised numbers (one, two, hundred, etc.), days of the week (Monday, etc.), months (January, etc.), abbreviations of days and months (Mon, Jun, etc.), ordinal numbers (first, second, etc.), cardinal prefixes (uni, bi, tri, etc.), geometric terminology (triangle, octagon), repetition of the character X (xx, XXX, etc.), among others (null, void, single, unity, decimal, etc.). We translated the list of words corresponding to numbers, months, and days of the week into French, Spanish, Italian, German, and Portuguese, using Google Translate with manual corrections [^4] For example, the word May was wrongly translated in Spanish to Puede (can) instead of Mayo.. Table 2 contains the full list of 445 words. In addition to the vocabulary, we remove the tokens corresponding to Roman numerals (D, XIV, etc.) and century names (XIXe, etc.). From this list of words, we create a list of 432 ignored tokens among the 32k tokens of the Llama 2 tokenizer. We ignore case, separation tokens, and plurals. Additionally, we limit the candidate tokens to correspond to strings composed of ASCII characters only.
@@
Category
Words
Digits
0, 1, 2, 3, 4, 5, 6, 7, 8, 9
Verbalised numbers
@@
Zero, One, Two, Three, Four, Five, Six, Seven, Eight, Nine, Ten, Eleven, Twelve, Thirteen, Fourteen, Fifteen, Sixteen, Seventeen, Eighteen, Nineteen, Twenty, Thirty, Forty, Fifty, Sixty, Seventy, Eighty, Ninety, Hundred, Thousand, Million, Billion, Trillion
Days of the week
@@
Monday, Tuesday, Wednesday, Thursday, Friday, Saturday, Sunday
Months
@@
January, February, March, April, May, June, July, August, September, October, November, December
Abbreviations of days and months
Jan, Feb, Mar, Apr, May, Jun, Jul, Aug, Sep, Oct, Nov, Dec, Mon, Tue, Wed, Thu, Fri, Sat, Sun
Numbers, months, and days in French
@@
Zéro, Un, Deux, Trois, Quatre, Cinq, Six, Sept, Huit, Neuf, Dix, Onze, Douze, Treize, Quatorze, Quinze, Seize, Dix-sept, Dix-huit, Dix-neuf, Vingt, Trente, Quarante, Cinquante, Soixante, Soixante-dix, Quatre-vingts, Quatre-vingt-dix, Cent, Mille, Million, Milliard, Janvier, Février, Mars, Avril, Mai, Juin, Juillet, Août, Septembre, Octobre, Novembre, Décembre, Lundi, Mardi, Mercredi, Jeudi, Vendredi, Samedi, Dimanche
Numbers, months, and days in Spanish
@@
Cero, Uno, Dos, Tres, cuatro, Cinco, Seis, Siete, Ocho, Nueve, Diez, Once, Doce, Trece, Catorce, Quince, Dieciséis, Diecisiete, Dieciocho, Diecinueve, Veinte, Treinta, Cuarenta, Cincuenta, Sesenta, Setenta, Ochenta, Noventa, Centenar, Mil, Millón, Billón, Enero, Febrero, Marzo, Abril, Mayo, Junio, Julio, Agosto, Septiembre, Octubre, Noviembre, Diciembre, Lunes, Martes, Miércoles, Jueves, Viernes, Sábado, Domingo
Numbers, months, and days in Italian
@@
Zero, Uno, Due, Tre, quattro, Cinque, Sei, Sette, Otto, Nove, Dieci, Undici, Dodici, Tredici, Quattordici, Quindici, Sedici, Diciassette, Diciotto, Diciannove, Venti, Trenta, Quaranta, Cinquanta, Sessanta, Settanta, Ottanta, Novanta, Centinaio, centi, Mille, milli, Milioni, Miliardi, Trilioni, Gennaio, Febbraio, Marzo, aprile, Maggio, Giugno, Luglio, agosto, settembre, ottobre, novembre, Dicembre, Lunedi, Martedì, Mercoledì, Giovedì, Venerdì, Sabato, Domenica
Numbers, months, and days in German
@@
Null, Eins, Zwei, Drei, Vier, Fünf, Sechs, Sieben, Acht, Neun, Zehn, Elf, Zwölf, Dreizehn, Vierzehn, Fünfzehn, Sechzehn, Siebzehn, Achtzehn, Neunzehn, Zwanzig, Dreißig, Vierzig, Fünfzig, Sechzig, Siebzig, Achtzig, Neunzig, Hundert, Tausend, Million, Milliarde, Billion, Januar, Februar, Marsch, April, Mai, Juni, Juli, August, September, Oktober, November, Dezember, Montag, Dienstag, Mittwoch, Donnerstag, Freitag, Samstag, Sonntag
Numbers, months, and days in Portuguese
@@
Zero, Um, Dois, Três, Quatro, Cinco, Seis, Sete, Oito, Nove, Dez, Onze, Doze, Treze, Quatorze, Quinze, Dezesseis, Dezessete, Dezoito, Dezenove, Vinte, Trinta, Quarenta, Cinquenta, Sessenta, Setenta, Oitenta, Noventa, Centenas, Mil, Milhão, Bilhão, Trilhão, Janeiro, Fevereiro, Marchar, abril, Maio, Junho, Julho, Agosto, Setembro, Outubro, novembro, dezembro, Segunda-feira, Terça-feira, Quarta-feira, Quinta-feira, Sexta-feira, Sábado, Domingo
Ordinal numbers
@@
First, Second, Third, Fourth, Fifth, Sixth, Seventh, Eighth, Ninth, Tenth, Eleventh, Twelfth, Thirteenth, Fourteenth, Fifteenth, Sixteenth, Seventeenth, Eighteenth, Nineteenth, Twentieth, Thirtieth, Fortieth, Fiftieth, Sixtieth, Seventieth, Eightieth, Ninetieth, Hundredth
Cardinal prefixes
@@
Uni, Bi, Tri, Quadri, Tetra, Penta, Quint, Sex, Hepta, Sept, Octa, Octo, oct, Nona, dec, Ennea
Geometric terminology
Triangle, Square, Hexagon, Pentagon
Repetition of the character X
@@
xx, xxx, xxxx, xxxxx, xxxxxx, xxxxxxx, xxxxxxxx, xxxxxxxxx, xxxxxxxxxx
Other
@@
Null, None, Void, Single, Singleton, Unity, Unique, Solo, Primary, Double, Pair, Twins, Duo, Binary, Secondo, Secondary, Seconda, Seconde, Couple, Twice, Handful, Triple, Trio, Triad, Quadruple, Quadr, Quartet, Quintet, Quintuple, Half-dozen, Sextet, Hexa, Septet, Heptagon, Septa, Octagon, Octet, Octave, Octopus, Nonagon, Nonet, Decimal, Dozen, Millionen
@@
Table 3 lists the system prompts used for the evaluation of suffixes in Section 5.3. “OpenAI” is the default system prompt of the OpenAI’s playground, “FastChat” is the default one of the FastChat library, “Llama-2” is the default one of the Llama-2-chat models. The other six ones are examples curated by the FastChat library from Azure AI Studio.
@@
Name
System Prompt
Source
Default Llama-2
@@
You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive in nature. If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information
Llama-2
Default OpenAI
You are a helpful assistant.
OpenAI’s playground
Default FastChat
@@
A chat between a curious human and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the human’s questions.
FastChat library
Shakespeare writing assistant
@@
You are a Shakespearean writing assistant who speaks in a Shakespearean style. You help people come up with creative ideas and content like stories, poems, and songs that use Shakespearean style of writing style, including words like "thou" and "hath". Here are some example of Shakespeare’s style: - Romeo, Romeo! Wherefore art thou Romeo? - Love looks not with the eyes, but with the mind; and therefore is winged Cupid painted blind. - Shall I compare thee to a summer’s day? Thou art more lovely and more temperate.
Azure AI Studio
IRS tax assistant

### v2

@@ Abstract
Large Language Model (LLM) services and models often come with legal rules on who can use them and how they must use them. Assessing the compliance of the released LLMs is crucial, as these rules protect the interests of the LLM contributor and prevent misuse. In this context, we describe the novel fingerprinting problem of Black-box Identity Verification (BBIV). The goal is to determine whether a third-party application uses a certain LLM through its chat function. We propose a method called Targeted Random Adversarial Prompt (TRAP) that identifies the specific LLM in use. We repurpose adversarial suffixes, originally proposed for jailbreaking, to get a pre-defined answer from the target LLM, while other models give random answers. TRAP detects the target LLMs with over 95% true positive rate at under 0.2% false positive rate even after a single interaction. TRAP remains effective even if the LLM has minor changes that do not significantly alter the original function.
@@ 1 Introduction
The recent proliferation of Large Language Models (LLMs) has drawn attention to several practical issues, such as model leaks, malicious usages and potential breaches of model licences. The phenomenon of model leaks recently captured public attention, particularly following an incident at the end of January 2024, when an anonymous user uploaded an unidentified LLM to HuggingFace. [^1] https://huggingface.co/miqudev/miqu-1-70b The CEO of Mistral subsequently confirmed that this was an internal model, leaked by an employee of an early access customer. [^2] https://twitter.com/arthurmensch/status/1752737462663684344 This event underscores the growing threat of internal breaches that LLM providers must contend with. LLM providers are also facing malicious usage of their technologies. For example, Yang and Menczer (2023) uncovered a network of social media bots utilizing ChatGPT to disseminate deceptive content. These bots promote suspicious websites and spread harmful comments, which violate OpenAI’s usage policies (OpenAI, 2024). Such challenges extend beyond proprietary LLMs to affect open-source models as well. A concrete example is Meta’s Llama 2 licence (Touvron et al., 2023), which forbids deceptive usage. Open-source LLM providers implement additional restrictions on model distribution. Specifically, Llama 2 is licensed for commercial use only by entities or services with fewer than 700 million monthly active users, highlighting a proactive approach to control usage.
@@
  • An analysis demonstrating TRAP’s reliability in identifying a model, even when other models are trained on the same data.
@@ 4.2 Fingerprints of Answers to Closed-Ended Questions
Even when the model outputs a unique fingerprint, the model does not reliably output it. Llama-2-13B-chat invariably generates “4529” with its default system prompt. Yet, it generates different answers when we change the system prompt. [^3] With a fixed prompt, Llama-2-13B-chat has a deterministic behaviour: it always outputs the same number. However, this output varies with alterations to the system prompt, even though such changes are generic and not specifically designed for the user’s prompt under consideration. Llama-2-13B-chat outputs “4289” with the OpenAI, Fastchat, and Xbox system prompts, “8273” with the marketing one, “4567” with the Json one, an “2341” with the Shakespeare one (see Appendix A for the system prompts). Guanaco-7B exhibited similar variability, underscoring the unreliability of this fingerprinting approach.
@@ 5.1 Targeted Random Adversarial Prompt (TRAP)
We apply the greedy coordinate gradient (GCG) in Zou et al. (2023). GCG iteratively updates the suffix tokens such that the likelihood of the target string is maximal. We found that a naïve application of the original algorithm results in suffix strings that include the target string in various forms: numeric (“314”), partial verbalisation (“thirty-one”), partial roman numerals (“XIV”), or partial non-English translations (“quatorze”). We have thus applied a filtering algorithm against all numeric strings [0-9] and verbalised numerals (see details in Sections 5.4 and A). The filtering is performed at each iteration.
@@
Difference with GCG.
Contrary to GCG (Zou et al., 2023), TRAP suffixes are not universal. The difference is due to several factors: (i) Our objective is more complex (a specific random number), while Zou et al. (2023)’s objective is “sure, here is” without constraints on what follows. (ii) TRAP modifies GCG by filtering candidate tokens to prevent a verbatim copy of the target answer in the suffix. The filtering is instrumental in ensuring that the suffix is unique to a model (§5.4). (iii) Zou et al. (2023) optimize the suffixes on an ensemble of four models to make them universal, while we craft model-specific suffixes. (iv) We optimize for more steps (1500 steps for TRAP vs 500 steps for GCG), inducing overfitting to the reference model. Zou et al. (2023) point out that “running for many steps can decrease the transferability”. For these three reasons, our suffixes are not universal (§5.2).
@@ 5.2 Trap Solves the BBIV Problem
Model Answer Length True Positive $\uparrow$ Invalid Answer False Positive $\downarrow$ Avg. Loss Avg. Best Step
Llama-2-7B-chat 3 99.3 % (991/998) 0.20 % (2/1000) 0.83 % 0.070 1244
4 95.2 % (940/987) 1.30 % (13/1000) 0.20 % 0.069 1172
5 76.8 % (751/978) 2.20 % (22/1000) 0.001 % 0.140 1221
Guanaco-7B 3 100 % (1000/1000) 0.00 % (0/1000) 0.81 % 0.157 1261
4 100 % (1000/1000) 0.00 % (0/1000) 0.01 % 0.155 1260
5 96.0 % (960/1000) 0.00 % (0/1000) 0.001 % 0.193 1234
Vicuna-7B 3 96.0 % (960/1000) 0.00 % (0/1000) 0.10 % 0.120 1218
4 97.0 % (970/1000) 0.00 % (0/1000) 0.01 % 0.133 1235
5 75.5 % (740/980) 2.00 % (20/1000) 0.001 % 0.210 1185
@@
Differentiating similar models.
TRAP can distinguish two models trained on the same dataset. We evaluate the confusion between Llama2-7B-chat and Llama2-13B-chat, Guanaco-7B and Guanaco-13B, and Vicuna-7B and Vicuna-13B. When a suffix is successful on the 7B model, the 13B model never outputs the target answer (Section B.2). TRAP identifies the reference model, even if other models are similar.
@@ 5.3 Robustness Analysis
Table 2: Ablation study. TRAP filters tokens to lower the false positive rate. False positive rate of 100 suffixes optimized on the model in row, using the token filtering in column. Answers are 4-digits long. In %.
Token Filtering
Optimized on TRAP Digits only None (GCG)
Llama2-7B-chat 0.13 1.35 3.82
Guanaco-7B 0.00 0.62 0.96
Vicuna-7B 0.00 1.41 0.83
5.4 Ablation Study
[…]
@@ 6 Conclusion and Discussion
While TRAP shows promise, it faces potential vulnerabilities to advanced countermeasures from third parties. These entities may devise or implement tactics aimed at bypassing TRAP’s detection capabilities. For example, a perplexity filter on the input could defend against TRAP. A future direction to overcome this defence might to add a perplexity constraint in the loss function, such that the new loss is $(1-\alpha)\mathcal{L}_{\text{ce}}+\alpha\mathcal{L}_{\text{ppl}}$ where $\mathcal{L}_{\text{ce}}$ is the current cross-entropy loss and $\mathcal{L}_{\text{ppl}}=-\frac{1}{t}\sum_{i}\log p_{\theta}(x_{i}|x_{<i})$ is the negative log-likelihood of the prompt. Jain et al. (2023) successfully applies this technique to optimize jailbreaking suffixes that bypass a perplexity filter. Despite our robustness analysis illustrated in Section 5.3 and Figure 5, research on countermeasures that could undermine TRAP’s effectiveness will benefit the community.
@@
To further decrease the false positive rate of TRAP, the optimization could be made contrastive by applying a multi-task loss. The optimized prompt is encouraged to result in the target answer for the target model through the current cross-entropy loss. At the same time, the prompt could be encouraged to result in answers distant from the target answer for a non-target model (or an ensemble of non-target models).
Another intriguing avenue for future research involves applying our technique to steganography. Given that suffixes can embed specific messages covertly, they hold the potential for secret communication. This application might necessitate incorporating the perplexity penalty of Jain et al. (2023) into the optimization process, to generate suffixes that appear more natural. Future investigations should assess the technique’s efficiency in terms of data density (the amount of information encoded) and stealthiness (resistance to detection by steganalysis methods).
@@ Acknowledgements
This work was supported by the NAVER corporation.
@@
For transparency and reproducibility, our code is available on GitHub at https://github.com/parameterlab/trap.
@@ Appendix A Experimental Settings
In our experiments, we optimized the suffixes on the Llama2-7B-chat, Guanaco-7B and the Vicuna-7B v1.3 models. Both models are aligned and fined-tuned for chat. Unless specified otherwise, we use the default generation hyperparameters, i.e., nucleus sampling with top-p of 0.9 and temperature of 0.6 for Llama-7B-chat model.
@@
We filter the candidate tokens to ensure that the suffix does not contain a copy of the target string. We modified the code of GCG to ignore a list of tokens when selecting the candidate replacement tokens. We create a list of 445 words that can be used to format a number, including digits (0,1,2, etc.), verbalised numbers (one, two, hundred, etc.), days of the week (Monday, etc.), months (January, etc.), abbreviations of days and months (Mon, Jun, etc.), ordinal numbers (first, second, etc.), cardinal prefixes (uni, bi, tri, etc.), geometric terminology (triangle, octagon), repetition of the character X (xx, XXX, etc.), among others (null, void, single, unity, decimal, etc.). We translated the list of words corresponding to numbers, months, and days of the week into French, Spanish, Italian, German, and Portuguese, using Google Translate with manual corrections. [^4] For example, the word May was wrongly translated in Spanish to Puede (can) instead of Mayo. Table LABEL:tab:data-words-filtered contains the full list of 445 words. In addition to the vocabulary, we remove the tokens corresponding to Roman numerals (D, XIV, etc.) and century names (XIXe, etc.). From this list of words, we create a list of 432 ignored tokens among the 32k tokens of the Llama 2 tokenizer. We ignore case, separation tokens, and plurals. Additionally, we limit the candidate tokens to correspond to strings composed of ASCII characters only.
@@
Category Words
Digits 0, 1, 2, 3, 4, 5, 6, 7, 8, 9
@@
Verbalised numbers Zero, One, Two, Three, Four, Five, Six, Seven, Eight, Nine, Ten, Eleven, Twelve, Thirteen, Fourteen, Fifteen, Sixteen, Seventeen, Eighteen, Nineteen, Twenty, Thirty, Forty, Fifty, Sixty, Seventy, Eighty, Ninety, Hundred, Thousand, Million, Billion, Trillion
@@
Days of the week Monday, Tuesday, Wednesday, Thursday, Friday, Saturday, Sunday
@@
Months January, February, March, April, May, June, July, August, September, October, November, December
Abbreviations of days and months Jan, Feb, Mar, Apr, May, Jun, Jul, Aug, Sep, Oct, Nov, Dec, Mon, Tue, Wed, Thu, Fri, Sat, Sun
@@
Numbers, months, and days in French Zéro, Un, Deux, Trois, Quatre, Cinq, Six, Sept, Huit, Neuf, Dix, Onze, Douze, Treize, Quatorze, Quinze, Seize, Dix-sept, Dix-huit, Dix-neuf, Vingt, Trente, Quarante, Cinquante, Soixante, Soixante-dix, Quatre-vingts, Quatre-vingt-dix, Cent, Mille, Million, Milliard, Janvier, Février, Mars, Avril, Mai, Juin, Juillet, Août, Septembre, Octobre, Novembre, Décembre, Lundi, Mardi, Mercredi, Jeudi, Vendredi, Samedi, Dimanche
@@
Numbers, months, and days in Spanish Cero, Uno, Dos, Tres, cuatro, Cinco, Seis, Siete, Ocho, Nueve, Diez, Once, Doce, Trece, Catorce, Quince, Dieciséis, Diecisiete, Dieciocho, Diecinueve, Veinte, Treinta, Cuarenta, Cincuenta, Sesenta, Setenta, Ochenta, Noventa, Centenar, Mil, Millón, Billón, Enero, Febrero, Marzo, Abril, Mayo, Junio, Julio, Agosto, Septiembre, Octubre, Noviembre, Diciembre, Lunes, Martes, Miércoles, Jueves, Viernes, Sábado, Domingo
@@
Numbers, months, and days in Italian Zero, Uno, Due, Tre, quattro, Cinque, Sei, Sette, Otto, Nove, Dieci, Undici, Dodici, Tredici, Quattordici, Quindici, Sedici, Diciassette, Diciotto, Diciannove, Venti, Trenta, Quaranta, Cinquanta, Sessanta, Settanta, Ottanta, Novanta, Centinaio, centi, Mille, milli, Milioni, Miliardi, Trilioni, Gennaio, Febbraio, Marzo, aprile, Maggio, Giugno, Luglio, agosto, settembre, ottobre, novembre, Dicembre, Lunedi, Martedì, Mercoledì, Giovedì, Venerdì, Sabato, Domenica
@@
Numbers, months, and days in German Null, Eins, Zwei, Drei, Vier, Fünf, Sechs, Sieben, Acht, Neun, Zehn, Elf, Zwölf, Dreizehn, Vierzehn, Fünfzehn, Sechzehn, Siebzehn, Achtzehn, Neunzehn, Zwanzig, Dreißig, Vierzig, Fünfzig, Sechzig, Siebzig, Achtzig, Neunzig, Hundert, Tausend, Million, Milliarde, Billion, Januar, Februar, Marsch, April, Mai, Juni, Juli, August, September, Oktober, November, Dezember, Montag, Dienstag, Mittwoch, Donnerstag, Freitag, Samstag, Sonntag
@@
Numbers, months, and days in Portuguese Zero, Um, Dois, Três, Quatro, Cinco, Seis, Sete, Oito, Nove, Dez, Onze, Doze, Treze, Quatorze, Quinze, Dezesseis, Dezessete, Dezoito, Dezenove, Vinte, Trinta, Quarenta, Cinquenta, Sessenta, Setenta, Oitenta, Noventa, Centenas, Mil, Milhão, Bilhão, Trilhão, Janeiro, Fevereiro, Marchar, abril, Maio, Junho, Julho, Agosto, Setembro, Outubro, novembro, dezembro, Segunda-feira, Terça-feira, Quarta-feira, Quinta-feira, Sexta-feira, Sábado, Domingo
@@
Ordinal numbers First, Second, Third, Fourth, Fifth, Sixth, Seventh, Eighth, Ninth, Tenth, Eleventh, Twelfth, Thirteenth, Fourteenth, Fifteenth, Sixteenth, Seventeenth, Eighteenth, Nineteenth, Twentieth, Thirtieth, Fortieth, Fiftieth, Sixtieth, Seventieth, Eightieth, Ninetieth, Hundredth
@@
Cardinal prefixes Uni, Bi, Tri, Quadri, Tetra, Penta, Quint, Sex, Hepta, Sept, Octa, Octo, oct, Nona, dec, Ennea
Geometric terminology Triangle, Square, Hexagon, Pentagon
@@
Repetition of the character X xx, xxx, xxxx, xxxxx, xxxxxx, xxxxxxx, xxxxxxxx, xxxxxxxxx, xxxxxxxxxx
@@
Other Null, None, Void, Single, Singleton, Unity, Unique, Solo, Primary, Double, Pair, Twins, Duo, Binary, Secondo, Secondary, Seconda, Seconde, Couple, Twice, Handful, Triple, Trio, Triad, Quadruple, Quadr, Quartet, Quintet, Quintuple, Half-dozen, Sextet, Hexa, Septet, Heptagon, Septa, Octagon, Octet, Octave, Octopus, Nonagon, Nonet, Decimal, Dozen, Millionen
@@
Table LABEL:tab:data-system-prompts lists the system prompts used for the evaluation of suffixes in Section 5.3. “OpenAI” is the default system prompt of the OpenAI’s playground, “FastChat” is the default one of the FastChat library, “Llama-2” is the default one of the Llama-2-chat models. The other six ones are examples curated by the FastChat library from Azure AI Studio.
@@
Name System Prompt Source
@@
Default Llama-2 You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive in nature. If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information Llama-2
Default OpenAI You are a helpful assistant. OpenAI’s playground
@@
Default FastChat A chat between a curious human and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the human’s questions. FastChat library
@@
Shakespeare writing assistant You are a Shakespearean writing assistant who speaks in a Shakespearean style. You help people come up with creative ideas and content like stories, poems, and songs that use Shakespearean style of writing style, including words like "thou" and "hath". Here are some example of Shakespeare’s style: - Romeo, Romeo! Wherefore art thou Romeo? - Love looks not with the eyes, but with the mind; and therefore is winged Cupid painted blind. - Shall I compare thee to a summer’s day? Thou art more lovely and more temperate. Azure AI Studio